OpenStudy (nathalyn):

Help 13.75=0.25b+0.50

OpenStudy (nathalyn):

@campbell_st

OpenStudy (anonymous):

Simplifying 13.75 = 0.25b + 0.50 Reorder the terms: 13.75 = 0.50 + 0.25b Solving 13.75 = 0.50 + 0.25b Solving for variable 'b'. Move all terms containing b to the left, all other terms to the right. Add '-0.25b' to each side of the equation. 13.75 + -0.25b = 0.50 + 0.25b + -0.25b Combine like terms: 0.25b + -0.25b = 0.00 13.75 + -0.25b = 0.50 + 0.00 13.75 + -0.25b = 0.50 Add '-13.75' to each side of the equation. 13.75 + -13.75 + -0.25b = 0.50 + -13.75 Combine like terms: 13.75 + -13.75 = 0.00 0.00 + -0.25b = 0.50 + -13.75 -0.25b = 0.50 + -13.75 Combine like terms: 0.50 + -13.75 = -13.25 -0.25b = -13.25 Divide each side by '-0.25'. b = 53 Simplifying b = 53

OpenStudy (nathalyn):

um thanks but i needed like this 3−18t+4=−26 3+−18t+4=−26 (−18t)+(3+4)=−26(Combine Like Terms) −18t+7=−26 −18t+7=−26 Subtract 7 from both sides. −18t+7−7=−26−7 −18t=−33 Divide both sides by -18 -18/-18=t -33/-18=? you cant divide -33/-18 so it would be a fraction you end up with t=11/6 t=11/6 is the final answer

OpenStudy (anonymous):

Oh ok